(a) If a 550-mL ordinary glass beaker is filled to the brim with ethyl alcohol at a temperature of 3.50°C, how much will overflow when their temperature reaches 22.0°C? The coefficient of volume expansion for ethyl alcohol is 1100 ✕ 10−6/°C. Assume the change in volume of the glass beaker is negligible compared to that of the liquid. mL

(b) How much less water would overflow under the same conditions? The coefficient of volume expansion for water is 210 ✕ 10−6/°C. Assume the change in volume of the glass beaker is negligible compared to that of the liquid. mL
Answer #1

a) delta V = beta * V0* delta T

= 1100 * 10-6 * 550 * (22 - 3.5)

= 11.1 mL

b) delta V = V0* alpha * delta T

= (1100 * 10-6 - 210 * 10-6)* 550 * (22 - 3.5)

= 9.0 mL